# Infinitely many genus two and three curves of small fixed positive rank

### v2

@@ Abstract.
We prove that there exist infinitely many genus two curves over $\mathbb{Q}$ whose Jacobians have ranks equal to a fixed number between 00 and $11$, and genus three curves over $\mathbb{Q}$ whose Jacobians have ranks equal to a fixed number between 00 and $4$. Furthermore, we consider a generalisation of these results over number fields. We prove that for any number field $K$, there are infinitely many genus two curves over $K$ whose Jacobians have ranks over $K$ equal to each of 0, 1, or 2.
@@ 1. Introduction
  • One part of the proof of the Hilbert 10th problem over the ring of integers $\mathcal{O}_{K}$, where $K$ is a number field, amounts to prove that for a given elliptic curve $E/\mathcal{O}_{K}$ of rank exactly 1, there exists its quadratic twist of rank also exactly 1, as proven recently for most number fields in [KP24]. One can also construct infinitely many elliptic curves of rank exactly 1 which are not $\overline{K}$-isomorphic, as proven independently in [Zyw25a].
@@
In the case of modular curves, using Eichler-Shimura relations, it is possible to decompose their Jacobians and compute their ranks by computing the ranks of their components. Still, this technique is not possible for most of the curves. Even for hyperelliptic curves, for which we can try to apply the 2-descent, it is not clear if the upper bound obtained from descent will match the number of $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. Knowing the existence of additional $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$ may not be helpful, because the 2-Selmer group of $\mathrm{Jac}(C)$ can be of large dimension, which is not sufficient to compute the rank of $\mathrm{Jac}(C)(K)$.
@@
For example, in [Gaj22], the first named author presents infinite families of curves over $\mathbb{Q}$ whose Jacobians have large ranks. Namely if we denote by $g(C)$ the genus of $C$, then an infinite family of hyperelliptic curves such that $\mathrm{Rank}(J(\mathbb{Q}))\geq g(C)-1$ is given in [Gaj22, Theorem 12(1)], whereas $\mathrm{Rank}(J(\mathbb{Q}))\geq g(C)$ is given in [Gaj22, Theorem 12(2)]. The main idea was to find curves with too many rational points that would violate Coleman’s bound [Col85], or Stoll’s improvement [Sto06]. However, giving an exact rank of such curves seems extremely difficult.
@@
In this manuscript, we shift our focus to studying arithmetic statistics of ranks of $\mathrm{Jac}(C)(K)$, where $C$ varies over an infinite family of curves. There is a wealth of previous studies which analyse the statistics of ranks over infinite family of curves, for example over universal families of elliptic curves [BS15], quadratic twist families of elliptic curves [KMR14, Smi22a, Smi22b, Smi25], and cubic twist families of elliptic curves [ABS22, KS24].
1.1. Our result and strategy
@@
The goal of this manuscript is to construct infinitely many genus two and three curves of fixed positive rank.
@@
In Section 2, we consider bielliptic genus two curves $C$ such that $\mathrm{Jac}(C)$ is isogenous to product of two elliptic curves $E_{1}$ and $E_{2}$. We divide the construction into 3 cases. In §2.1, we consider the case over number fields $K$ where both $E_{1}$ and $E_{2}$ do not have any $K$-rational $2$-torsion points. In §2.2, we examine elliptic curves over $\mathbb{Q}$ having marked rational $2$-torsion points. In §2.3, we use elliptic curves over $\mathbb{Q}$ having all rational $2$-torsion points. We then construct families of genus two curves $C$ by fixing one curve $E_{2}$ of the prescribed rank $r$, and quadratically twisting $E_{1}$ except in §2.4, which we explain later. This allows us to utilise results from previous literature (such as [Nag29, Fre84, KMR14]) that there are infinitely many such twists of rank 0 or 1 depending on the choice of $E_{1}$. We can hence construct infinitely many bielliptic curves of prescribed rank. For $K=\mathbb{Q}$, we construct the desired families for $0\leq r\leq 11$.
The strategy we take for genus three curves is similar. We consider bielliptic curves $C$ such that $\mathrm{Jac}(C)$ is isogenous to a product of an elliptic curve $E$ and the Jacobian of a genus two curve $C^{\prime}$. We then construct families of genus three curves $C$ by fixing the genus two curve $C$ of prescribed rank $r$, and quadratically twisting $E$. We then use previous studies [Kan13, Smi22a] to obtain that over $K=\mathbb{Q}$, there are infinitely many such twists of rank 0. This allows us to construct infinitely many bielliptic curves of ranks 00, $1$, and $2$ over $\mathbb{Q}$.
@@
For most of our constructions, it is crucial to fix an elliptic curve of the prescribed rank and to aim for twists of rank 0. Other strategies, for instance, to fix an elliptic curve of rank 0 and to aim for infinitely many twists of rank $r$, seem much more difficult, as in principle, it is much easier to show the existence of infinite families of curves of rank 0 rather than positive rank. We prove that there are not many rational points on these curves, as explained in §4.1. We use positive rank twists in one of the examples in §2.2.
@@
In §2.4, we also construct infinitely many curves over any number field $K$ of genus two, whose rank of the Jacobian over $K$ is each of 0, 1, and 2. These constructions use groundbreaking results in [Smi22b, KP25], which give an existence of infinitely many quadratic twists of a certain elliptic curve over $K$ of rank 0 [Smi22b] or 1 [KP25]. The conditions for elliptic curves for which their theorems can be applied are more complicated than those over $\mathbb{Q}$. Unlike the previous three subsections, the non-zero rank examples we construct crucially depend on rank 1 quadratic twist families. Checking whether the conditions appearing in [KP25] are applicable to our construction of genus two curves is an additional technicality we address in this subsection. The heart of our result for number fields lies in Theorem 2.8, where we propose an intricate construction for such genus two curves. Our work can be considered as a continuation of the recent progress by pioneering constructions of infinitely many elliptic curves of rank exactly equal to $1$ over any number field [KP25, Zyw25a]. More concretely, our work is an application of the results from [KP25] and of the same question for rank 0 [Smi22b].
We finally ask for an existence of infinitely many genus two or three curves defined over $\mathbb{Q}$ with a fixed positive rank of its Jacobian over $\mathbb{Q}$, but with an additional condition that the Jacobian is absolutely simple. This condition makes it a difficult question, see §4.2.
@@ 1.2. Explicit families of positive rank
Even though there are several instances of families of curves of fixed positive rank, such as [ABS22, KP25, Zyw25a], not so many give examples of families with explicit parametrisation. There are nevertheless some previous studies which give explicit parametrisation of families of elliptic curves over $\mathbb{Q}$ with rank $1$, such as those proven by Monsky [Mon92] and Dasgupta and Voight [DV18], those with rank $2$ as proven recently by Zywina [Zyw25b]. We give explicit examples of curves of genus two and a fixed rank between 0 and 4 in §2.2 and §2.3.
@@ 2. Genus two curves
We consider bielliptic curves $C$ of genus two; then $\mathrm{Jac}(C)\sim E_{1}\times E_{2}$. In our constructions we fix one curve, say $E_{1}$, and we vary $E_{2}$. Note that, in general, $E_{1}\times E_{2}$ is not isogenous to a Jacobian of a genus two curve, but in our cases, the curves $E_{1}$ and $E_{2}$ are “compatible” so it is. More formally, we can glue them so that one gets a bielliptic curve of genus 2, whose Jacobian is isogenous to $E_{1}\times E_{2}$. We provide the constructions for all possible rational two torsion subgroups when considered over $\mathbb{Q}$, i.e., $E_{1}[2](\mathbb{Q})\cong E_{2}[2](\mathbb{Q})\in\{\langle O\rangle,\mathbb{Z}/2\mathbb{Z},\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\}$. However, our construction in §2.1 works over more general number fields.
@@ 2.1. No rational two-torsion
$\begin{split}C_{D}&\colon y^{2}=x^{6}+D,\\ E_{1,D}&\colon y^{2}=x^{3}+D,\\ E_{2,D}&\colon y^{2}=x^{3}+D^{2}.\end{split}$ (1)
Then there exists an isogeny $\phi\colon\mathrm{Jac}(C_{D})\to E_{1,D}\times E_{2,D}$. A direct application of [BD11, Theorem 3.2] shows that $C_{D}$ is isogenous to a product of two elliptic curves whose Weierstrass equations are $E^{\prime}_{1,D}\colon y^{2}=x^{3}+D$ and $E^{\prime}_{2,D}\colon y^{2}=x^{4}+Dx$. A change of variables given by $X:=\frac{D}{x}$ and $Y:=\frac{Dy}{x^{2}}$ gives the Weierstrass equation for $E_{2,D}$.
@@
[…]
Then for any $D\in\mathcal{D}_{a}$, the Weierstrass equation for $E_{2,D}$ can be rewritten as $E\colon y^{2}=x^{3}+a^{2}$, which is independent of the choice of $D$. Because $K(E[2])/K$ is an $S_{3}$ extension, $K(E_{1,D}[2])/K$ is also an $S_{3}$ extension. We note that $\{E_{1,D}\}_{D\in D_{a}}$ is a family of quadratic twists of the elliptic curve $E^{\prime}\colon y^{2}=x^{3}+a$. By [KMR14, Theorem A], there exists an infinite subset $\mathcal{D}^{0}_{a}\subset\mathcal{D}_{a}$ such that $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(E_{1,D}/K)=0$, hence $\mathrm{Rank}E_{1,D}(K)=0$, for every $D\in\mathcal{D}^{0}_{a}$.
We note that for any $D\in\mathcal{D}_{a}$ there are infinitely many $D^{\prime}\in\mathcal{D}_{a}$ such that $C_{D}\not\cong C_{D^{\prime}}$. Indeed, suppose one chooses $D^{\prime}$ such that $DD^{\prime}\not\in\mathcal{O}_{K}^{2}$ and the number of prime ideals of $K$ dividing $D^{\prime}$ that are coprime to $(6a)\subset\mathcal{O}_{K}$ is strictly greater than that of $D$. Then there exists a place of $K$ such that $C_{D^{\prime}}$ has bad reduction whereas $C_{D}$ does not. Hence, we have that $\{C_{D}\}_{D\in D^{0}_{a}}$ is an infinite set of hyperelliptic curves such that $\mathrm{Rank}(\mathrm{Jac}(C_{D})(K))=r$. ∎
@@
For $0\leq r\leq 4$, we apply Theorem 2.1 for the following values of $a=2,3,15,427,13\cdot 19\cdot 23\cdot 43$ for $r=0,1,2,3,4$, respectively. The corresponding ranks can easily be checked in Magma [BCP97]. For $5\leq r\leq 11$, we apply the main result from [ER04]. We note that the elliptic curve $y^{2}=x^{3}-432k^{2}$ is $3$-isogenous to the elliptic curve $y^{2}=x^{3}+16k^{2}$ [CP09, Section 1]. We can then set $a=4k$ for the following values of $k$ to produce infinitely many genus two curves of rank $r$ as provided below.
$\begin{split}r=5&:k=3\cdot 7\cdot 11\cdot 13\cdot 163,\\ r=6&:k=3\cdot 73\cdot 103\cdot 439,\\ r=7&:k=3\cdot 13\cdot 19\cdot 41\cdot 139\cdot 271,\\ r=8&:k=2\cdot 3\cdot 5\cdot 7\cdot 11\cdot 13\cdot 17\cdot 29\cdot 41\cdot 47\cdot 59,\\ r=9&:k=2\cdot 5\cdot 37\cdot 41\cdot 53\cdot 73\cdot 1231\cdot 4831,\\ r=10&:k=2\cdot 3\cdot 5\cdot 7\cdot 23\cdot 31\cdot 37\cdot 43\cdot 83\cdot 109\cdot 151\cdot 421,\\ r=11&:k=3\cdot 5\cdot 7\cdot 13\cdot 19\cdot 23\cdot 31\cdot 43\cdot 59\cdot 61\cdot 73\cdot 79\cdot 103\cdot 109\cdot 157\cdot 457.\end{split}$ (3)
@@
The condition of $K(E[2])/K$ being an $S_{3}$ extension amounts to $\zeta_{3}\notin K$ (necessary condition) and $3]{a}\notin K$. Therefore, for all number fields $K$ such that $3\nmid[K:\mathbb{Q}]$ and $\zeta_{3}\notin K$, this condition is satisfied. Otherwise, for fields with $\zeta_{3}\notin K$, but with $3]{a}\in K$, one can try to find another elliptic curve $E^{\prime}\colon y^{2}=x^{3}+a^{\prime 2}$ satisfying the conditions of Theorem 2.1.
@@
Suppose $K=\mathbb{Q}$. Using the main results from [Smi22a, Smi22b, Smi25], we can nullify the condition from Theorem 2.1 that $K(E[2])/K$ is an $S_{3}$ extension. Let $a\in\mathbb{Q}^{\times}$ be a fixed number. As before, let $E:y^{2}=x^{3}+a^{2}$. Then at least $50\%$ of genus two curves in $\{C_{D}\}_{D\in\mathcal{D}_{a}}$ have Jacobians of rank equal to $\mathrm{Rank}(E(\mathbb{Q}))$, and $100\%$ of genus two curves in $\{C_{D}\}_{D\in\mathcal{D}_{a}}$ have Jacobians of rank at most $\mathrm{Rank}(E(\mathbb{Q}))+1$.

### v3

# Infinitely many hyperelliptic curves of small genus and small fixed rank, and of any genus and rank two

@@ Abstract.
We prove that for any number field $K$ and any fixed genus $g\geq 2$, there are infinitely many non-isomorphic hyperelliptic curves of genus $g$ over $K$ whose Jacobians have rank over $K$ equal to each of 0, 1, or 2. As an example of our method, over $\mathbb{Q}$, we prove that there exist infinitely many non-isomorphic hyperelliptic curves of genus two, whose Jacobians have rank equal to a fixed number between $1$ and $11$, genus three and four curves with rank between $1$ and $4$, and genus five and six with rank between $1$ and $3$.
@@ 1. Introduction
  • One part of the proof of the Hilbert 10th problem over the ring of integers $\mathcal{O}_{K}$, where $K$ is a number field, amounts to prove that there is an abelian variety $A/K$ and a quadratic extension $L/K$ such that $\mathrm{Rank}(A(L))=\mathrm{Rank}(A(K))>0$. This was first proven very recently in the groundbreaking work by Peter Koymans and Carlo Pagano [KP24] who constructed an elliptic curve $E/\mathcal{O}_{K}$ and a quadratic extension $L/K$ such that $\mathrm{Rank}(E(L))=\mathrm{Rank}(E(K))=1$. One can also construct infinitely many elliptic curves of rank exactly 1 which are not $\overline{K}$-isomorphic, as proven independently in [Zyw25b]. A different proof of Hilbert 10th problem using positive rank of Jacobians of some twist families of hyperelliptic curves was shown by another groundbreaking work by Levent Alpöge, Manjul Bhargava, Wei Ho, and Ari Shnidman [ABHS25].
@@
In the case of modular curves, using Eichler-Shimura relations, it is possible to decompose their Jacobians and compute their rank by computing the rank of their components. Still, this technique is not possible for most of the curves. Even for hyperelliptic curves, for which we can try to apply the 2-descent, it is not clear if the upper bound obtained from descent will match the number of $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. Knowing the existence of additional $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$ may not be helpful, because the 2-Selmer group of $\mathrm{Jac}(C)$ can be of large dimension, which is not sufficient to compute the rank of $\mathrm{Jac}(C)(K)$.
@@
For example, in [Gaj22], the first named author presents infinite families of curves over $\mathbb{Q}$ whose Jacobians have large rank. Namely if we denote by $g(C)$ the genus of $C$, then an infinite family of hyperelliptic curves such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))\geq g(C)-1$ is given in [Gaj22, Theorem 12(1)], whereas $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))\geq g(C)$ is given in [Gaj22, Theorem 12(2)]. The main idea was to find curves with too many rational points that would violate Coleman’s bound [Col85], or Stoll’s improvement [Sto06]. However, giving an exact rank of such curves seems extremely difficult.
@@
In this manuscript, we shift our focus to studying arithmetic statistics of rank of $\mathrm{Jac}(C)(K)$, where $C$ varies over an infinite family of curves. There is a wealth of previous studies which analyse the statistics of rank over infinite family of curves, for example over universal families of elliptic curves [BS15], quadratic twist families of elliptic curves [KMR14, Smi22a, Smi22b, Smi25], cubic twist families of elliptic curves [ABS22, KS24], quartic twist families of elliptic curves over $\mathbb{Q}(i)$ [Sav], and infinitely many elliptic curves of fixed rank between 0 and 4 over any number field [Zyw25a].
1.1. Our results and strategy
@@
The goal of this manuscript is to construct infinitely many non-isomorphic higher genus curves of fixed positive rank.
Notation. Throughout this manuscript, we will not write explicitly the adjective non-isomorphic and the construction of infinitely many curves will always assume they are non-isomorphic. The proof that they are non-isomorphic in each case is straightforward, because we construct the curves by gluing twists of hyperelliptic curves, and for infinitely many twists, we introduce a new prime of bad reduction (using that we twist by square-free elements).
Motivation and difficulties. To the best of our knowledge, the highest exact rank of an elliptic curve over $\mathbb{Q}$ up to this date is 20, found by Elkies and Klagsburn [EK20] (whereas the same people found and elliptic curve of rank at least 29, which is the highest rank elliptic curve over $\mathbb{Q}$, but it is not known if 29 is the exact rank of this curve). Thus, in principle, we can construct an abelian variety over $\mathbb{Q}$ of dimension $g$ and rank $r=20g$, just take $E^{g}$, where $E$ is the known rank 20 curve.
However, there are great mathematical merits to restrict our attention to Jacobians of curves rather than just random abelian varieties because, as we explained before, the curves play a crucial role in investigating many of the most important mathematical questions, and to study and understand their arithmetic, the knowledge of the rank of its Jacobian is one of the most essential information one can obtain. In particular, even though we admit that it is more difficult to compute the rank of a simple Jacobian, as we also mentioned, studying the modular curves lead to so many exciting and groundbreaking results, and they, in principle, have decomposable Jacobians. Hence, it also motivates studying decomposable Jacobians.
In this manuscript, we pay particular focus to studying Jacobians of hyperelliptic curves. Recall that for $g\geq 2$, the dimensions of the moduli space of principally polarised abelian varieties of dimension $g$, the moduli space of curves of genus $g$, and the moduli space of hyperelliptic curves of genus $g$ over $\mathbb{C}$ are respectively given by $\frac{g(g+1)}{2}$, 3g33g-3, and 2g12g-1. For $g\geq 3$, it is clear that the hyperelliptic locus is a thin subset of both moduli spaces of principally polarised abelian varieties and moduli spaces of genus $g$ curves, so the aforementioned strategy of simply taking products of lower dimensional abelian varieties does not construct Jacobians of hyperelliptic curves. Even for $g=2$, it is not clear whether a product of some two elliptic curves is isogenous to the Jacobian of a hyperelliptic curve over a given number field $K$. These considerations highlight one of the subtleties that this manuscript aims to address: we provide explicit gluing procedure for two hyperelliptic curves to produce a new hyperelliptic curve of higher genusdefined over the given number field $K$, thus focusing our problem to thin subset of moduli space of principally polarised abelian varieties.
Another subtlety lies in ensuring that the gluing procedure allows us to incorporate many deep and technical results regarding constructions of hyperelliptic curves whose Jacobians have small fixed rank. Our article demonstrates that all relevant conditions appearing in such previous works can be checked and satisfied even after undertaking the gluing procedure for two hyperelliptic curves over $K$. In particular, we find examples of curves such that effectively, when gluing them, we can consider one of them as a fixed curve (in fact, we use curves which are isomorphic to that one), and then we show that we can apply the known results about the twists of the other one, or we allow twisting both curves with a good control of the rank of both twists concurrently.
Organisation. Observant readers may realise that Section 2 and Section 3 contain similar arguments regarding gluing two hyperelliptic curves to construct a new hyperelliptic curve whose Jacobian has fixed positive rank. The reason why we introduce the genus two case first before the higher genus case in this manuscript is to present our method to potential readers in a more intuitive and accessible manner. Our contributions to this project initially started with genus two hyperelliptic curves. The gluing method is simpler and easier to state for genus 2 hyperelliptic curves, and we hope introducing these techniques first will allow readers to feasibly follow through the generalisations to cases for higher genus hyperelliptic curves.
@@
In Section 2, we consider bielliptic genus two curves $C$ such that $\mathrm{Jac}(C)$ is isogenous to product of two elliptic curves $E_{1}$ and $E_{2}$. We divide the construction into 3 cases. In §2.1, we consider the case over number fields $K$ where both $E_{1}$ and $E_{2}$ do not have any $K$-rational 22-torsion points. In §2.2, we examine elliptic curves over $\mathbb{Q}$ having marked rational $2$-torsion points. In §2.3, we use elliptic curves over $\mathbb{Q}$ having all rational $2$-torsion points. We then construct families of genus two curves $C$ by fixing one curve $E_{2}$ of the prescribed rank $r$, and quadratically twisting $E_{1}$ except in §2.5, which we explain later. This allows us to utilise results from previous literature (such as [Nag29, Fre84, KMR14]) that there are infinitely many such twists of rank 0 or 1 depending on the choice of $E_{1}$. We can hence construct infinitely many bielliptic curves of prescribed rank. For $\boldsymbol{K=\mathbb{Q}}$, we construct the desired families for $\boldsymbol{0\leq r\leq 11}$.
@@
For most of our constructions, it is crucial to fix an elliptic curve of the prescribed rank and to aim for twists of rank 0. Other strategies, for instance, to fix an elliptic curve of rank 0 and to aim for infinitely many twists of rank $r$, seem much more difficult, as in principle, it is much easier to show the existence of infinite families of curves of rank 0 rather than positive rank. We prove that there are not many rational points on these curves, as explained in §2.4. We use positive rank twists in one of the examples in §2.2 and §2.5.
@@
In §2.5, we also construct infinitely many curves over any number field $K$ of genus two, whose rank of the Jacobian over $K$ is each of 0, 1, and 2. These constructions use groundbreaking results in [Smi22b, KP25], which give an existence of infinitely many quadratic twists of a certain elliptic curve over $K$ of rank 0 [Smi22b] or 1 [KP25]. The conditions for elliptic curves for which their theorems can be applied are more complicated than those over $\mathbb{Q}$. Unlike the previous three subsections, the non-zero rank examples we construct crucially depend on rank 1 quadratic twist families. Checking whether the conditions appearing in [KP25] are applicable to our construction of genus two curves is an additional technicality we address in this subsection. The heart of our result for number fields lies in Theorem 2.8, where we propose an intricate construction for such genus two curves. Our work can be considered as a continuation of the recent progress by pioneering constructions of infinitely many elliptic curves of rank exactly equal to $1$ over any number field [KP25, Zyw25b]. More concretely, our work is an application of the results from [KP25] and of the same question for rank 0 [Smi22b].
Generalizing our strategy explored for genus two curves, in Section 3 we consider hyperellitpic curves $C$ of genus $g\geq 3$ such that $\mathrm{Jac}(C)$ is isogenous to product of two Jacobians of hyperelliptic curves $C_{1}$ and $C_{2}$. In §3.1, we construct infinitely many curves over any number field $K$ of any genus at least $3$, whose rank of the Jacobian over $K$ is each of 0, 1, and 2. The overall strategy of the proof remains identical, but we employ different previous results to generate such curves. We still use [Yu19, BM, Smi22b] for the existence of infinitely many quadratic twists of Jacobians of a certain hyperelliptic curve over $K$ of rank 0. But we use the recent breakthrough by Peter Koymans and Adam Morgan [KM] for the existence of infinitely many such quadratic twists of rank $1$. We check in §3.1 that the conditions appearing in [Smi22b] and [KM] are once again applicable to our construction of higher genus curves, as we highlight in Theorem 3.10 and Theorem 3.13. In remaining subsections, we give examples of infinitely many hyperelliptic curves over $\mathbb{Q}$ of genus 3, 4, 5, and 6 with fixed positive rank. To the best of our knowledge, our work is the first attempt to utilize gluing method in the setting of generating infinitely many higher genus hyperelliptic curves with positive fixed rank.
The groundbreaking work by Peter Koymans and Adam Morgan [KM] in fact answers the previously posed challenge problem, which asks for generating infinitely many absolutely simple Jacobians of hyperelliptic curves over $\mathbb{Q}$ with rank $1$. We mention the formulation of the previously posed question in Section 4.
@@ 1.2. Explicit families of positive rank
Even though there are several instances of families of curves of fixed positive rank, such as [ABS22, KP25, Zyw25b], not so many give examples of families with explicit parametrisation. There are nevertheless some previous studies which give explicit parametrisation of families of elliptic curves over $\mathbb{Q}$ with rank $1$, such as those proven by Monsky [Mon92] and Dasgupta and Voight [DV18]. Also, very recently, the question of whether there are infinitely many rank 2 elliptic curves over $\mathbb{Q}$ was proved by Zywina [Zyw25c], who, impressively, constructs elliptic curves non-isomorphic over $\overline{\mathbb{Q}}$, and we can say that these results are semi-explicit. While the elliptic curves are given by concrete equations, these equations depend on a certain primality condition that can be viewed as non-explicit, but we still believe that this result deserves to be mentioned in this section. We give explicit examples of curves of genus two and a fixed rank between 0 and 4 in §2.2 and §2.3.
The Magma [BCP97] code for our computations is available at https://github.com/StevanGajovic/Ranks-Weil-Pairings.
@@ 2. Genus two curves
We consider bielliptic curves $C$ of genus two; then $\mathrm{Jac}(C)\sim E_{1}\times E_{2}$. In our constructions we fix one curve, say $E_{1}$, and we vary $E_{2}$. Recall that, in general, $E_{1}\times E_{2}$ is not isogenous to a Jacobian of a genus two curve over $K$, but in our cases, the curves $E_{1}$ and $E_{2}$ are “compatible” so it is. More formally, we can glue them so that one gets a bielliptic curve of genus 2, whose Jacobian is isogenous to $E_{1}\times E_{2}$. We provide the constructions for all possible rational two torsion subgroups when considered over $\mathbb{Q}$, i.e., $E_{1}[2](\mathbb{Q})\cong E_{2}[2](\mathbb{Q})\in\{\langle O\rangle,\mathbb{Z}/2\mathbb{Z},\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\}$. However, our construction in §2.1 works over more general number fields.
@@ 2.1. No rational two-torsion
(1) $\begin{split}C_{D}&\colon y^{2}=x^{6}+D,\\ E_{1,D}&\colon y^{2}=x^{3}+D,\\ E_{2,D}&\colon y^{2}=x^{3}+D^{2}.\end{split}$
Then there exists an isogeny $\phi\colon\mathrm{Jac}(C_{D})\to E_{1,D}\times E_{2,D}$. A direct application of [BD11, Theorem 3.2] shows that $\mathrm{Jac}(C_{D})$ is isogenous to a product of two elliptic curves whose Weierstrass equations are $E^{\prime}_{1,D}\colon y^{2}=x^{3}+D$ and $E^{\prime}_{2,D}\colon y^{2}=x^{4}+Dx$. A change of variables given by $X:=\frac{D}{x}$ and $Y:=\frac{Dy}{x^{2}}$ gives the Weierstrass equation for $E_{2,D}$.
@@
(2) $\mathcal{D}_{a}:=\left\{a\cdot m^{3}\in\mathcal{O}_{K}\colon m\in\mathcal{O}_{K}^{\times}\text{ is square-free}\right\}.$
Then for any $D\in\mathcal{D}_{a}$, the Weierstrass equation for $E_{2,D}$ can be rewritten as $E\colon y^{2}=x^{3}+a^{2}$, which is independent of the choice of $D$. Because $K(E[2])/K$ is an $S_{3}$ extension, $K(E_{1,D}[2])/K$ is also an $S_{3}$ extension. We note that $\{E_{1,D}\}_{D\in D_{a}}$ is a family of quadratic twists of the elliptic curve $E^{\prime}\colon y^{2}=x^{3}+a$. By [KMR14, Theorem A], there exists an infinite subset $\mathcal{D}^{0}_{a}\subset\mathcal{D}_{a}$ such that $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(E_{1,D}/K)=0$, hence $\mathrm{Rank}E_{1,D}(K)=0$, for every $D\in\mathcal{D}^{0}_{a}$. Hence, we have that $\{C_{D}\}_{D\in D^{0}_{a}}$ is an infinite set of hyperelliptic curves such that $\mathrm{Rank}(\mathrm{Jac}(C_{D})(K))=r$. ∎
@@
For $0\leq r\leq 4$, we apply Theorem 2.1 for the following values of $a=2,3,15,427,13\cdot 19\cdot 23\cdot 43$ for $r=0,1,2,3,4$, respectively. The corresponding rank can easily be checked in Magma [BCP97]. For $5\leq r\leq 11$, we apply the main result from [ER04]. We note that the elliptic curve $y^{2}=x^{3}-432k^{2}$ is $3$-isogenous to the elliptic curve $y^{2}=x^{3}+16k^{2}$ [CP09, Section 1]. We can then set $a=4k$ for the following values of $k$ to produce infinitely many genus two curves of rank $r$ as provided below.
(3) $\begin{split}r=5&:k=3\cdot 7\cdot 11\cdot 13\cdot 163,\\ r=6&:k=3\cdot 73\cdot 103\cdot 439,\\ r=7&:k=3\cdot 13\cdot 19\cdot 41\cdot 139\cdot 271,\\ r=8&:k=2\cdot 3\cdot 5\cdot 7\cdot 11\cdot 13\cdot 17\cdot 29\cdot 41\cdot 47\cdot 59,\\ r=9&:k=2\cdot 5\cdot 37\cdot 41\cdot 53\cdot 73\cdot 1231\cdot 4831,\\ r=10&:k=2\cdot 3\cdot 5\cdot 7\cdot 23\cdot 31\cdot 37\cdot 43\cdot 83\cdot 109\cdot 151\cdot 421,\\ r=11&:k=3\cdot 5\cdot 7\cdot 13\cdot 19\cdot 23\cdot 31\cdot 43\cdot 59\cdot 61\cdot 73\cdot 79\cdot 103\cdot 109\cdot 157\cdot 457.\end{split}$
@@
The condition of $K(E[2])/K$ being an $S_{3}$ extension amounts to $\zeta_{3}\notin K$ (necessary condition) and $\sqrt[3]{a}\notin K$. Therefore, for all number fields $K$ such that $3\nmid[K:\mathbb{Q}]$ and $\zeta_{3}\notin K$, this condition is satisfied. Otherwise, for fields with $\zeta_{3}\notin K$, but with $\sqrt[3]{a}\in K$, one can try to find another elliptic curve $E^{\prime}\colon y^{2}=x^{3}+a^{\prime 2}$ satisfying the conditions of Theorem 2.1.
@@
Suppose $K=\mathbb{Q}$. Using the main results from [Smi22a, Smi22b, Smi25], we can nullify the condition from Theorem 2.1 that $K(E[2])/K$ is an $S_{3}$ extension by allowing $a$ to be a cube. Let $a\in\mathbb{Q}^{\times}$ be a fixed number. As before, let $E:y^{2}=x^{3}+a^{2}$. Then at least $50\%$ of genus two curves in $\{C_{D}\}_{D\in\mathcal{D}_{a}}$ have Jacobians of rank equal to $\mathrm{Rank}(E(\mathbb{Q}))$, and $100\%$ of genus two curves in $\{C_{D}\}_{D\in\mathcal{D}_{a}}$ have Jacobians of rank at most $\mathrm{Rank}(E(\mathbb{Q}))+1$.